OpenStudy (imamermaid29):

Will fan & Medal Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y A. (3,-1/2) B. (0,-2) C. (0, -11/2) D. (7,-5/2)

OpenStudy (anonymous):

do you know what the substitution method is?

OpenStudy (imamermaid29):

no... i mean maybe if i saw it, math is my really bad subject

OpenStudy (anonymous):

you take the answers (a,b, etc) then put for example (0,-2) one of the number sentences would be 0 – 2 = – 2 * -2.

OpenStudy (anonymous):

\[x – 2 = – 2y\] can be written as \(x=-2y+2\) you can replace \(x\) in \(4x+2y=11\) by \(-2y+2\) and solve \[4(-2y+2)+2y=11\]

OpenStudy (anonymous):

that is the "substitution" method, where you substitute an expression in one variable for another

OpenStudy (anonymous):

now the question is, do you know how to solve \[4(-2y+2)+2y=11\] for \(y\)?